# On Variational Perspectives To The Graph Isomorphism Problem

### v2

@@ V.2 VQE Results
Figure 3: The illustration of the energy for VQE (y-axis) for pairs of graphs with four nodes in respect with number of iterations (x-axis). The colors represent different runs on a class of graphs. It can be seen that after 400400 iterations the energies cluster.
VI Conclusion and Future Work
We have tested two different quantum algorithmic approaches, namely the QAOA and VQE, towards obtaining solutions of small case-instances for the Graph Isomorphism Problem, where the Hamiltonian is formulated using the QUBO formulation. We have shown through numerical simulations that the results of both algorithms support each other in identifying isomorphic and non-isomorphic pairs for four, and five nodes in a non-noisy device. Thus, we have proposed a new application area for variational quantum algorithms.
@@ VI Conclusion and Future Work
We have demonstrated some empirical evidence that agree with previous literature Szegedy (2019) in regards to QAOA energies of isomorphic graphs clustering together. Additionally, we have also shown that similar phenomena happen for VQE simulations as well. A further extension of the investigation pursued in this work involves expanding the idea of energies clustering for families of isomorphic graphs and running the formulation on current devices with greater qubit number. Another natural next step for this line of research is to scale the variational formulations to graphs with 6–8 nodes using more efficient encodings that reduce qubit overhead. While the quadratic growth of qubits in the QUBO formulation presents a bottleneck, recent work on compact encodings and symmetry exploitation could make moderately larger instances tractable. In parallel, testing these formulations on real quantum hardware such as IBM’s superconducting processors or Google’s Sycamore device would provide valuable insights into the effects of noise, decoherence, and limited circuit depth. Employing error mitigation strategies, including zero-noise extrapolation, measurement error correction, and dynamical decoupling, would allow us to benchmark the robustness of the observed energy clustering under realistic device conditions and thereby bridge the gap between proof-of-principle simulations and experimental feasibility.

### v3

@@ VI.2 VQE as a variational probe of the isomorphism Hamiltonian
Figure 5: VQE convergence behavior for four-node graph pairs. Shown are variational energies as a function of optimization iteration for multiple independent VQE runs applied to the same four-node graph pair. Each curve corresponds to a distinct random initialization of the variational parameters (distinguished by both color and line style to aid accessibility). Panels (a) and (b) correspond to two representative graph classes considered in this work (specified in the main text). After approximately 400 iterations, all runs converge to a narrow band of final energies, illustrating energy clustering, i.e., low variance in converged energies across runs. This clustering indicates reproducible convergence to low-energy regions of the Hamiltonian landscape, but does not by itself imply discrimination between isomorphic and non-isomorphic graph instances.
To disentangle algorithmic effects from properties intrinsic to the Hamiltonian encoding, we perform a controlled structural ablation experiment in which graph topology is varied while graph isomorphism is held fixed. Starting from the five-node complete graph, edges are systematically removed to generate a sequence of increasingly sparse graphs, as shown in Fig. 4(a). For each graph in this sequence, we consider a pair of identical copies, thereby ensuring that all tested instances are strictly isomorphic by construction.
Figure 4(b) reports the final QAOA energies obtained for each such isomorphic pair. Because graph labeling is held constant up to permutation, any variation in the observed energies cannot be attributed to relabeling artifacts or optimization noise, but instead reflects changes in the underlying Hamiltonian induced by graph structure. The resulting energy profile demonstrates that the variational energy minimum shifts systematically as edges are removed, while remaining tightly clustered within each isomorphism class.
This figure therefore serves as a calibration experiment: it verifies that the Hamiltonian responds sensitively and smoothly to controlled structural changes in the input graph, independent of labeling. Importantly, the observed clustering should not be interpreted as a decision criterion for graph isomorphism. In particular, near-degenerate energies among distinct graph instances (e.g., graphs (d), (f), and (g) in Fig. 4) directly impact algorithmic performance by preventing reliable discrimination based on variational energy alone. Rather, the clustering illustrates how structural features of the graph are encoded into the energy landscape, providing a baseline against which the limitations of energy-based discrimination discussed later can be understood.
The observation of energy clustering is scientifically relevant because it probes the structure of the variational energy landscape induced by the graph isomorphism Hamiltonian. Clustering demonstrates that, for a fixed graph instance, the Hamiltonian admits broad low-energy basins that are consistently reached by variational optimization, independent of initialization. By energy clustering, we mean the empirical observation that repeated variational optimizations of the same Hamiltonian—initialized from different random parameters—converge to final energies whose variance is small compared to the total energy scale of the problem. In this sense, clustering refers to concentration of the final energies within a narrow band, rather than the emergence of well-separated energy gaps between different graph instances. This indicates that the Hamiltonian encoding is stable and reproducible, rather than dominated by optimizer noise or parameter sensitivity.
At the same time, the absence of well-separated clusters across different graph instances highlights a fundamental limitation: while variational algorithms reliably identify low-energy regions, these regions are not uniquely associated with valid isomorphisms or distinct graph classes. Consequently, energy clustering serves both as a diagnostic of Hamiltonian structure and as evidence that energy-based criteria alone are insufficient for scalable graph isomorphism discrimination.
Taken together, these results highlight the complementary roles of QAOA and VQE in analyzing the graph isomorphism Hamiltonian. QAOA, due to its shallow depth and limited expressivity, serves as a lightweight diagnostic tool for probing coarse features of the energy landscape and identifying structural trends. VQE, by contrast, provides a more faithful approximation of the Hamiltonian ground state, at the expense of substantially increased computational resources.
Crucially, neither approach circumvents the fundamental difficulty of enforcing global permutation constraints within a variational framework. As demonstrated elsewhere in this work, low variational energy does not reliably imply feasibility of the decoded permutation, particularly at shallow depth. Consequently, while energy clustering is a robust feature of the Hamiltonian encoding, it should be interpreted as a property of the variational energy landscape rather than as a complete or scalable solution to the graph isomorphism problem.
VII Conclusion and Future Work
In this work, we examined variational quantum approaches to the graph isomorphism problem through a detailed study of the energy landscapes induced by standard QUBO encodings. Using both QAOA and VQE, we analyzed small graph instances and demonstrated that variational optimization reliably produces structured, reproducible low-energy behavior for isomorphic graph pairs. Controlled ablation experiments further confirmed that these energy responses are driven by graph topology rather than labeling artifacts, validating the fidelity of the Hamiltonian encoding in the variational setting.
At the same time, our results make clear that low variational energy alone does not constitute a reliable or scalable criterion for graph isomorphism. We explicitly observed substantial overlap between the low-energy spectra of isomorphic and non-isomorphic instances, as well as frequent convergence to energetically favorable but infeasible configurations that violate permutation constraints. These effects persist even when incorporating trajectory-level information and classical post-processing, underscoring intrinsic limitations of energy-based discrimination at shallow circuit depth.
Taken together, these findings position variational algorithms not as solvers for the graph isomorphism problem, but as diagnostic tools for probing the structure and limitations of Hamiltonian encodings in the NISQ regime. By characterizing both the strengths and failure modes of QAOA and VQE in this context, this work provides a principled baseline for future efforts that seek to incorporate feasibility-preserving mixers, constraint-aware ansätze, or hybrid classical-quantum strategies to move beyond purely energetic heuristics.
@@ VII Conclusion and Future Work
A further extension of the investigation pursued in this work involves expanding the idea of energies clustering for families of isomorphic graphs and running the formulation on current devices with greater qubit number. Another natural next step for this line of research is to scale the variational formulations to graphs with 6–8 nodes using more efficient encodings that reduce qubit overhead. While the quadratic growth of qubits in the QUBO formulation presents a bottleneck, recent work on compact encodings and symmetry exploitation could make moderately larger instances tractable. In parallel, testing these formulations on real quantum hardware such as IBM’s superconducting processors or Google’s Sycamore device would provide valuable insights into the effects of noise, decoherence, and limited circuit depth. Employing error mitigation strategies, including zero-noise extrapolation, measurement error correction, and dynamical decoupling, would allow us to benchmark the robustness of the observed energy clustering under realistic device conditions and thereby bridge the gap between proof-of-principle simulations and experimental feasibility.